OpenStudy (anonymous):

Pleasee help!! I need the equation of a circle with a center at (0,0) and a radius of 5. Also, an equation of a circle with a center of (0.0) and a diameter of 20. And an equation of a circle with a center at (1,2) and a diameter of 15.

geerky42 (geerky42):

General equation of circle is \((x-h)^2 + (y-k)^2 = r^2\), where (h,k) is center of the circle and r is radius

OpenStudy (anonymous):

so the first answer would be 25?? @geerky42

OpenStudy (anonymous):

No, just subsitute (h,k) in the formula that geerky42 gave you

OpenStudy (anonymous):

i cant crack it

OpenStudy (anonymous):

but if h,k is (0,0)?

OpenStudy (anonymous):

So, just subistute and you'll see.

OpenStudy (anonymous):

2+2=5^2?

OpenStudy (anonymous):

\[(x-h)^{2}+(y-k)^{2}=r^{2} where, h=0, k=0, r=5\] Substitute thos values in the equation

OpenStudy (anonymous):

so (x-0)^2 +(y-0)^2=5?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

and that is my equation?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

that was really easy! THANK YOUU

OpenStudy (anonymous):

Now do the same thing with the other two problems

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

wait @charliealbert what if I the problem only provides the center and the diameter?

OpenStudy (anonymous):

What is the diameter?

OpenStudy (anonymous):

center (0,0) and diameter 20

OpenStudy (anonymous):